Based on the following balance sheet and income statement. Identify all financial ratios listed for each year. *Feel free to just do the first year! I need to check to make sure my answers are appropriate. Thanks!

ANNUAL BALANCE SHEET
($ Thousands)
31-Dec-15 31-Dec-16
31-Dec-15 31-Dec-16
ASSETS Profitability Ratios (in %)
Cash & Short-Term Investments $17.11 $25.05 Return on equity
Net Receivables $1.62 $1.84 Return on assets
Inventories $42.72 $39.68 Return on invested capital
Prepaid Expenses $0.00 $0.00 Profit margin
Other Current Assets $3.38 $2.93 Gross margin
Total Current Assets $64.84 $69.49 Turnover-Control Ratios
Asset turnover
Gross Plant, Property & Equipment $63.75 $58.28 Fixed-asset turnover
Accumulated Depreciation $36.16 $32.61 Inventory turnover
Collection period (days)
Net Plant, Property & Equipment $27.58 $25.67 Days' sales in cash
Intangibles $5.88 $6.01 Payables period
Other Assets $6.16 $5.81
Leverage and Liquidity Ratios
TOTAL ASSETS $104.47 $106.98 Assets to equity
Debt to assets
LIABILITIES Debt to equity
Long Term Debt Due In One Year $0.00 $0.00 Times interest earned
Accounts Payable $10.59 $11.91 Current ratio
Taxes Payable $1.87 $2.01 Acid test
Accrued Expenses $7.19 $6.91
Other Current Liabilities $1.88 $1.85
Total Current Liabilities $21.54 $22.67
Long Term Debt $6.95 $19.79
Deferred Taxes $1.16 $2.32
Other Liabilities $3.04 $2.43
TOTAL LIABILITIES $32.68 $47.21
EQUITY
Common Stock $0.07 $0.06
Capital Surplus $27.25 $24.31
Retained Earnings $73.89 $61.10
Less: Treasury Stock $29.42 $25.71
TOTAL EQUITY $71.79 $59.77
TOTAL LIABILITIES & EQUITY $104.47 $106.98
Common Shares Outstanding $5.14 $5.05
Annual Income Statement
($ Thousands, except per share )
Sales $179.24 $164.28
Cost of Goods Sold $95.71 $91.99
Gross Profit $83.53 $72.29
Selling, General, & Administrative Exp. $56.36 $50.65
Operating Income Before Deprec. $27.17 $21.64
Depreciation,Depletion,&Amortization $5.85 $5.89
Operating Profit $21.33 $15.74
Interest Expense $0.88 $0.56
Non-Operating Income/Expense $0.93 $0.31
Special Items $0.00 $0.00
Pretax Income $21.38 $15.49
Total Income Taxes $7.23 $5.60
Income Before Extraordinary
Items & Discontinued Operations $14.15 $9.90
Savings Due to Common Stock Equiv. $0.00 $0.00
Adjusted Net Income $14.15 $9.90
EPS Basic from Operations $0.26 $0.19
EPS Diluted from Operations $0.25 $0.19
Dividends Per Share $0.02 $0.00
Com Shares for Basic EPS $5.06 $5.12
Com Shares for Diluted EPS $5.21 $5.27
Answer #1
31-Dec-15
Profitability Ratios (in %) Formula Calculation Ratio
Return on equity Net income / Shareholder equity =14.15/71.79 19.71%
Return on assets Net income / Total assets =14.15/104.47 13.54%
Profit margin Net income / Sales =14.15/179.24 7.89%
Gross margin Gross profit / Sales =83.53/179.24 46.60%
Turnover-Control Ratios
Asset turnover Sales / Total assets =179.24/104.47 1.72
Fixed-asset turnover Sales / Fixed assets =179.24/27.58 6.50
Inventory turnover Sales / Inventory =179.24/42.72 4.20
Collection period (days) 365/ (Sales / Receivable) =365/(179.24/1.62) 3.30
Payables period 365/ (COGS / Payable) =365/(95.71/10.59) 40.39
Leverage and Liquidity Ratios
Assets to equity Total assets / Equity =104.47/71.79 1.46
Debt to assets (Long term debt + Short term debt) / Total assets =6.95/104.47 0.07
Debt to equity (Long term debt + Short term debt) / Equity =6.95/71.79 0.10
Times interest earned Operating profit / Interest expense =21.33/0.88 24.24
Current ratio Current assets / Current liabilities =64.84/21.54 3.01
Acid test (Current assets - Inventory)/ Current liabilities =(64.84-42.72)/21.54 1.03